OpenStudy (anonymous):

Sketch the graph of F(x)=2^x^-1

OpenStudy (anonymous):

F(x)=2^x^-1 = f(x) = 2^(1/x)

OpenStudy (anonymous):

Make an x-y chart, from -2 to 2 and plot

OpenStudy (anonymous):

*-1 to 1

OpenStudy (anonymous):

Well te part i dont get is when you get negative exponents .i have to give atleast 5 points on the graph

OpenStudy (anonymous):

2^-1 = (1/2)

OpenStudy (anonymous):

The number (2/1) is flipped to (1/2)

OpenStudy (anonymous):

So. Put it under 1 and then solve he exponent?

OpenStudy (anonymous):

No, you want to solve for y. Lets make x=-1 y=2^(1/(-1))

OpenStudy (anonymous):

when x=-1, y=(1/2)

OpenStudy (anonymous):

(1/-1) = -1. 2^-1 = y = (1/2)

OpenStudy (anonymous):

Now make x=0

OpenStudy (anonymous):

I got1/2

OpenStudy (anonymous):

Wha?

OpenStudy (anonymous):

2^(1/0) =2^infinity = undefined

OpenStudy (anonymous):

you cannot divide by 0

OpenStudy (anonymous):

2^(1/1) = 2^1 = 2

OpenStudy (anonymous):

So we got (-1,0.5) and (1,2)

OpenStudy (anonymous):

If i put 1 in the equation, i get 2^0 whic is 1

OpenStudy (anonymous):

No, 2^(1/x). (1/1) = 1 ; 2^1 =2

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=y%3D2^x^-1%3B The graph should look like this

OpenStudy (anonymous):

Thin makes nonsense

OpenStudy (anonymous):

No sense

OpenStudy (anonymous):

Are you sure you typed the equation correctly?

OpenStudy (anonymous):

Please go slow

OpenStudy (anonymous):

It's not 2^(x-1) is it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I wish i could draw it

OpenStudy (anonymous):

@ajprincess

OpenStudy (anonymous):

You there?

OpenStudy (ajprincess):

yup I am.

OpenStudy (ajprincess):

x^(-1) can be written as 1/x